# Learned frequency-domain scattered wavefield solutions using neural operators

### v1

@@ I Introduction
With the development of modern machine learning techniques comes the possibility of using neural networks to surrogate numerical modeling. Physics-informed neural networks (PINNs) have recently been employed for frequency-domain wavefield modeling, demonstrating their potential in solving complex wave equations in irregular geometry through meshless solutions [2, 3, 4]. Despite their multi-frequency and multi-source capabilities [5], PINNs’ need for retraining or transfer learning for new velocities limits their generalization. Consequently, developing neural networks for a range of parametric PDEs becomes crucial. Neural operators [6] have shown great potential in solving for parametric PDEs, which provide instant PDE solutions given various coefficients and boundary conditions. Recently, they have also been used for seismic wavefield simulation, e.g., 2D time-domain wavefield simulation [7]. However, learning seismic waveform modeling in the time domain faces several challenges: (1) Learning the mapping directly from input 2D velocities and source locations to 3D wavefield volumes (spanning 2D spatial and 1D time-lapse domains) requires huge memory cost [8]; (2) Learning the mapping between the snapshots at different times within recurrent composition though decreases the memory cost, will be subject to error propagation [9] as well as the relatively huge inference cost due to the iterative nature. Thus, learning in the frequency domain using neural operators, which avoids the above limitations, could efficiently handle parametric wave equations.
Although current neural operator implementations for the frequency domain wavefield solutions offer memory efficiency, they typically often use the end-to-end framework to learn the mapping from the velocities, source locations, and frequencies to the corresponding wavefield solutions without additional feature engineering [8, 10]. Specifically, they treat the source locations as binary masks and frequencies as an additional channel for inputting neural operators with constant values. However, without considering the physics of seismic wavefields, this approach offers suboptimal inputs for convolution-based feature extraction networks. Thus, it requires additional modifications to the network architecture, e.g., paralleled Fourier neural operator [10] or limits the target frequencies of modeling [8] to make the neural operators handle various velocities and source configurations (e.g., locations and frequency) simultaneously.
@@ II-A Operator learning for Partial differential equations
where $a\in\mathcal{A}$ denotes the parameters of the operator $\mathcal{L}$, such as coefficient functions, and $u\in\mathcal{U}$ is the physical phenomenon we aim to solve for and belonging to solution space $\mathcal{U}$. The parameters and the solution are a function of location $x$ defined within domain $\Omega$. For learning an operator, we assume that, for any $a\in\mathcal{A}$, there exists a unique solution $u=\mathcal{F}(a)\in\mathcal{U}$ satisfying Eq. (1), then $\mathcal{F}$ is the solution operator of Eq. (1). Regarding the acoustic wave equation in seismic modeling, $a$ includes the source locations, frequencies, and velocity models, and $u$ represents the corresponding simulated wavefield.
@@ II-C Learning for the scattered wavefield
Following the conventional supervised learning paradigm, we randomly sample source locations, frequencies, and velocities $\{v^{i},x_{s}^{i},z_{s}^{i},\omega^{i}\}$ to compute the background wavefield $U_{0}^{i}$ by solving Equation 3, and then generate the corresponding wavefield $\delta U^{i}$ by solving Equation 4 numerically for each sample. Using these samples, we train a neural network $f_{\theta}$ with input $(v^{i},U_{0}^{i})$ to predict the scattered wavefield (or the full wavefield by adding a residual connection from the input background wavefield directly to output) using a mean squared error loss
@@
where $N$ is the number of training samples. As for the choice of the neural network, we select the commonly used benchmark network, the Fourier neural operator (FNO) [6], to test the framework. The full pipeline of the approach is shown in Figure 1.
@@ III-A Synthetic tests on OpenFWI dataset
Here, we first applied the proposed method to the OpenFWI dataset, focusing on the ’CurveVelA’ velocity class [12], as illustrated in Figure 3. This class of velocities includes curved layers, and the value of velocity within the layers gradually increases with depth. This test involved a comprehensive training set comprising 9,000 samples, encompassing various velocities, source locations, and frequencies, and a corresponding validating set comprising 1,000 samples. Although the Fourier neural operator can learn the mapping between infinite function space, the training data here are discretized samples with a spatial resolution of 139$\times$139. The training curve and validation error analysis are shown in Figure 2, and as we can see, the training converges well.
Figure 2: The training history for the ’CurveVelA’ test. Top: the training loss curve; Bottom: the validation error, which is measured by the relative L2 norm between the predictions and reference simulated results.
@@
III-B More realistic velocity models
To further test the effectiveness of the proposed method, we choose more realistic velocity models to test, e.g., the combinations of Overthrust, Marmousi models, and so on. Figure 5 shows samples of velocities in the training dataset. The number of training samples is 5,400, while that for validation is 600. Figure 6 shows the training loss curve and the validation error at each epoch. We also train FNO up to 1000 epochs, and it converges well. We similarly obtain the numerical reference results on unseen cases (unseen configurations of velocities, frequencies, and source locations) with a spatial interval of 0.025 km in both $x$ and $z$ directions and compare them with the predictions from the FNO. Figure 7 shows a comparison between the numerical solutions and the network predictions. Like the OpenFWI dataset tests, the FNO predicts accurate wavefields instantly.
@@ III-B More realistic velocity models
Figure 6: The training history for the realistic models test. Top: the training loss curve; Bottom: the validation error, which is measured by the relative L2 norm between the predictions and reference simulated results.
@@
III-C Generalization test with reference frequency:
@@ IV Discussions
The proposed method has shown good performance on the OpenFWI and realistic velocity models. It frees us at inference time from numerical simulations, in which the computational complexity is high. Taking the test on realistic models as an example, the numerical simulation, executed in parallel on an Intel(R) Xeon(R) Gold 6230R CPU @ 2.10GHz, for each velocity and frequency at a resolution of 256$\times$256 requires 6.4 seconds, and that at a resolution of 512$\times$512 requires 38.3 seconds. Using our method, the elapsed time on an NVIDIA A100 GPU (we do not fully utilize the GPU due to the relatively small input size and batch size) is 0.036 seconds for the resolution of 256$\times$256, 1.1 seconds for the resolution of 512$\times$512. Those speedups come from two main aspects: (1) we avoid repeated simulations; (2) we can easily utilize the latest hardware and software advances of accelerated computing [17] because we build the algorithm on machine learning. However, we recognize the challenges associated with the preparation of data-label pairs required for training the neural operator, particularly as the resolution of training samples increases. Additionally, the generalization capability to unseen velocities hinges on the diversity and volume of the training samples. While these challenges can be mitigated to some extent through physics-constrained learning, our focus in this paper is on introducing the neural operator with the background wavefield as input. Furthermore, the generalization prowess of our method across varied velocity models holds promising implications for applications in full waveform inversion (FWI), potentially eliminating the need for retraining and achieving real-time FWI. In our implementation, we decided to have the actual velocity as input and the scattered wavefield as output. However, the method can be equally applied by alternatively having the perturbation velocity (difference between the true and homogeneous background) as input and also the option of having the scattered or the full wavefield as output. As mentioned earlier, the full wavefield can also be obtained by employing a residual connection between the input background wavefield and the scattered wavefield.

### v2

@@ I Introduction
With the development of modern machine learning techniques comes the possibility of using neural networks to surrogate numerical modeling. Physics-informed neural networks (PINNs) have recently been employed for time-domain and frequency-domain wavefield modeling, demonstrating their potential in solving complex wave equations in irregular geometry through meshless solutions [2, 3, 4, 5, 6, 7, 8]. Despite their multi-source and multi-frequency capabilities [9], PINNs’ need for retraining or transfer learning for new velocities limits their generalization. Consequently, developing neural networks for a range of parametric PDEs becomes crucial. Neural operators [10] have shown great potential in solving for parametric PDEs, which provide instant PDE solutions given various coefficients and boundary conditions. Recently, they have also been used for seismic wavefield simulation, e.g., 2D/3D time-domain wavefield simulation [11, 12, 13]. However, learning seismic waveform modeling in the time domain faces several challenges: (1) Learning the mapping directly from input 2D velocities and source locations to 3D wavefield volumes (spanning 2D spatial and 1D time-lapse domains) requires huge memory cost [14]; (2) Learning the mapping between the snapshots at different times within recurrent composition though decreases the memory cost, will be subject to error propagation [15] as well as the relatively huge inference cost due to the iterative nature. Thus, learning in the frequency domain using neural operators, which avoids the above limitations, could efficiently handle parametric wave equations.
Although current neural operator implementations for the frequency domain wavefield solutions offer memory efficiency, they typically often use the end-to-end framework to learn the mapping from the velocities, source locations, and frequencies to the corresponding wavefield solutions without additional feature engineering [14, 16, 17]. Specifically, they treat the source locations as binary masks and frequencies as an additional channel for inputting neural operators with constant values. However, without considering the physics of seismic wavefields, this approach offers suboptimal inputs for convolution-based feature extraction networks. Thus, it requires additional modifications to the network architecture, e.g., paralleled Fourier neural operator [17] or limits the target frequencies of modeling [14] to make the neural operators handle various velocities and source configurations (e.g., locations and frequency) simultaneously.
@@
To summarize, the main contributions of our paper are as follows:
  • We propose to utilize analytical background wavefields to implicitly represent the source location and frequency information for neural operator-based seismic wavefield-based simulation, yielding better accuracy and generalization.
  • We use a reference frequency strategy to enhance the performance of the trained neural operator on a larger unseen simulation domain.
  • We validate the proposed method on a class of velocity models from OpenFWI and more realistic models and demonstrate the effectiveness of the proposed method.
@@ II-A Operator learning for Partial differential equations
where $a\in\mathcal{A}$ denotes the parameters of the operator $\mathcal{L}$, such as coefficient functions, and $u\in\mathcal{U}$ is the physical phenomenon we aim to solve for and belonging to solution space $\mathcal{U}$. The parameters and the solution are a function of location $x$ defined within domain $\Omega$. For learning an operator, we assume that, for any $a\in\mathcal{A}$, there exists a unique solution $u=\mathcal{F}(a)\in\mathcal{U}$ satisfying (1), then $\mathcal{F}$ is the solution operator of (1). Regarding the acoustic wave equation in seismic modeling, $a$ includes the source locations, frequencies, and velocity models, and $u$ represents the corresponding simulated wavefield.
@@ II-C Learning for the scattered wavefield
Following the conventional supervised learning paradigm, we randomly sample source locations, frequencies, and velocities $\{v^{i},x_{s}^{i},z_{s}^{i},\omega^{i}\}$ to compute the background wavefield $U_{0}^{i}$ by solving (3), and then generate the corresponding wavefield $\delta U^{i}$ by solving (4) numerically, with a optimal 9-point finite difference method, for each sample. Using these samples, we train a neural network $f_{\theta}$ with input $(v^{i},U_{0}^{i})$ to predict the scattered wavefield (or the full wavefield by adding a residual connection from the input background wavefield directly to output) using a mean squared error loss
@@
where $N$ is the number of training samples. As for the choice of the neural network, among many neural networks for parametric PDEs such as ResNet [22], Unet [23], UNO [24], LordNet [15], graph neural networks [25], transformers [26] and so on, we select the commonly used benchmark network, the Fourier neural operator (FNO) [10], to test the framework. The full pipeline of the approach is shown in Figure 1.
@@ III-A Synthetic tests on OpenFWI dataset
Here, we first applied the proposed method to the OpenFWI dataset, focusing on the ’CurveVelA’ velocity class [19], as illustrated in Figure 3. This class of velocities includes curved layers, and the value of velocity within the layers gradually increases with depth. This test involved a comprehensive training set comprising 9,000 samples, encompassing various velocities, source locations, and frequencies, and a corresponding validating set comprising 1,000 samples. The validation error is estimated using the relative L2 norm error, which equals the Euclidean distance from the prediction to the ground truth divided over the Euclidean norm of the ground truth. Although the Fourier neural operator can learn the mapping between infinite function space, the training data here are discretized samples with a spatial resolution of 139$\times$139. The training curve and validation error analysis are shown in Figure 2, and as we can see, the training converges well.
Figure 2: The training history for the ’CurveVelA’ test. Top: the training loss curve; Bottom: the validation error.
@@
III-B Comparison with conventional neural operator-based methods
As mentioned above, conventional neural operator-based methods treat the source locations as binary masks and frequencies as additional channels. This approach offers suboptimal inputs for a convolution-based feature extraction network, yielding a degraded performance. In this subsection, we demonstrate this weakness of the conventional methods, especially compared to our proposed method for enhancing the wavefield representation. For a fair comparison, we use the same training and validation datasets, as well as the same configurations for the FNOs, to train the neural operator using the different approaches (pipelines). We specifically use three different approaches here: the conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), the conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), and our method. To have the evaluation metrics consistent and comparable, for the full wavefield output, the calculation of the relative L2 norm error is also done by the calculation of the L2 norm of the errors divided by the L2 norm of the scattered wavefield. The training loss curves and the validation error curves are shown in Figure 6. It is obvious that our approach, implicitly embedding the source location and frequency information into the background wavefield, can improve the convergence, as well as improve the accuracy and generalization ability a lot compared to conventional neural operator-based methods.
Figure 6: The comparison between different approaches on the training histories and validation errors for the real and imaginary parts.
We further compare their performance on the unseen velocities with 200 samples. As shown in Table I, our method achieves the best accuracy compared to the conventional approaches with almost the same computational cost.
TABLE I: The training cost and accuracy comparison between different approaches
     Approaches      Relative L2 error      Training cost
     Real      Imag.      (hours)
     Conventional in full out      0.3290      0.3271      15.714
     Conventional in residual out      0.6935      0.6943      15.771
     Our method      0.2598      0.2599      15.759
III-C More realistic velocity models
To further test the effectiveness of the proposed method, we choose more realistic velocity models to test, e.g., the combinations of Overthrust, Marmousi models, and so on. Figure 5 shows samples of velocities in the training dataset. The number of training samples is 5,400, while that for validation is 600. Figure 7 shows the training loss curve and the validation error at each epoch. We also train FNO up to 1000 epochs, and it converges well. We similarly obtain the numerical reference results on unseen cases with a spatial interval of 0.025 km in both $x$ and $z$ directions and compare them with the predictions from the FNO. Figure 8 shows a comparison between the numerical solutions and the network predictions. Like the OpenFWI dataset tests, the FNO predicts accurate wavefields instantly.
Figure 7: The training history for the realistic models test. Top: the training loss curve; Bottom: the validation error.
@@ III-C More realistic velocity models
III-D Generalization test with reference frequency:
@@ IV Discussions
The proposed method has shown good performance on the OpenFWI and realistic velocity models. It frees us at inference time from numerical simulations, in which the computational complexity is high. Taking the test on realistic models as an example, the numerical simulation for each velocity and frequency, executed in parallel on an Intel(R) Xeon(R) Gold 6230R CPU @2.10GHz, at a resolution of 139$\times$139 requires 1.27 seconds, and that at a resolution of 256$\times$256 and 512$\times$512 require 2.79 and 8.64 seconds, respectively. Using our method, the elapsed time on an NVIDIA A100 GPU (we do not fully utilize the GPU due to the relatively small input size and batch size) is 0.008 seconds for the resolution of 139$\times$139, 0.012 seconds for the resolution of 256$\times$256, and 0.031 seconds for the resolution of 512$\times$512. The main computationally intensive part of the proposed method is the training. In our experiments, it took 15.75 hours for training the neural operator for the ”CurveVelA velocity models, and 29.25 hours for the realisitc velocity models. Noteworthy, all of the training are performed offline. When trying to apply our method, indeed the neural operator-based wavefield solutions, on the downstream tasks such as reverse time migration or full waveform inversion, which requires massive simulations on different frequencies and source locations, the speed-up ratio of our method compared to conventional numerical methods will significantly increase, which will easily justify the expensive training. Taking multiple frequencies simulations on a resolution of 139$\times$139 as an example, the elapsed times for learned wavefield solutions and numerical finite-difference method with respect to the number of frequencies are shown in Figure 13. Since the learned wavefield solutions make it easy to parallel the prediction without additional modification (like parallel computing required for the numerical method to handle multiple frequencies simultaneously), instead, we just increase the batch size of the input tensors. It is obvious that within 16 simulations, the elapsed time for the learned wavefield solutions did not exceed 1 s, while the elapsed time for the numerical simulation without additional modification increased linearly. Due to the saturation of computational resources and memory bandwidth, the cost of neural operator-based wavefield solutions is slightly increased, causing a non-linear increase in the speed-up ratio. In summary, at inference, it is much faster than the conventional numerical solvers. Those speedups come from two main aspects: (1) we avoid repeated simulations; (2) we can easily utilize the latest hardware and software advances of accelerated computing [29] because we build the algorithm on machine learning. However, we recognize the challenges associated with the preparation of data-label pairs required for training the neural operator, particularly as the resolution of training samples increases.
In terms of generalization, taking the neural operator trained on ”CurveVelA” as an example, we further evaluate its performance on other entirely different types of velocity models such as ”FlatVelA” (no curvature in the velocity model), ”FlatFaultA” (no curvature, but include faults), and ”CurveFaultA” (with curvature and faults). For each class, we test five velocity models; for each velocity model, we test eight frequencies; and for each frequency, we test three shots. We show the relative L2 norm error for each class in Figure 11. The errors for high-frequency prediction are large, while those for low-frequency are low. We show three good results and one poor case among these tests in Figure 12. For those testing velocities with quite different features compared to the training datasets, our method can still generalize well for the low-frequency wavefield predictions. However, it fails to accurately predict high-frequency wavefields on the ”CurveFaultA” velocity model. The generalization capability to unseen velocities, hinges on the diversity and volume of the training samples. While these challenges can be mitigated to some extent through physics-constrained learning, our focus in this paper is on introducing the neural operator with the background wavefield as input.
Figure 11: The relative L2 norm errors for each class of models. Regarding the sample index from 0 to 120, each of the 24 consecutive numbers represents a velocity model, and each of the three consecutive numbers represents a frequency, which increases in sequence. Each frequency value corresponds to three different shots.
Figure 12: The samples of the out-of-distribution test. The first three rows are the samples on ”FlatVelA”, ”FlatFaultA”, and ”CurveFaultA” for low-frequency simulation, while the fourth row is the worst case on ”CurveFaultA” for high-frequency simulation.
Furthermore, the generalization prowess of our method across varied velocity models holds promising implications for applications in full waveform inversion (FWI), potentially eliminating the need for retraining and achieving real-time FWI. In our implementation, we decided to have the actual velocity as input and the scattered wavefield as output. However, the method can be equally applied by alternatively having the perturbation velocity (difference between the true and homogeneous background) as input and also the option of having the scattered or the full wavefield as output. Here, we take CurveVelA as an example to compare the difference between our method and our method with full wavefield output, shown in Figure 14. Their performance is almost the same. However, as mentioned earlier, the full wavefield can also be obtained by employing a residual connection between the input background wavefield and the scattered wavefield.
Figure 13: The efficiency comparison in terms of elapsed time for the learned wavefield solutions and the numerical finite-difference method on a resolution of 139$\times$139 with respect to different number of frequencies.
Figure 14: The comparison between FNO with scattered wavefield as output and that with full wavefield as output on the validation errors for the real and imaginary parts
Finally, to further highlight the effectiveness of the proposed method on different neural network architectures, we compare the different approaches such as such as conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), our method with full wavefield outputs, and our method with scattered wavefield outputs on different neural network architectures. We pick ResNet [22] and Unet [23] for the study. The results are shown in Tables II and III. The results demonstrate that our method is generally effective with other neural networks.
TABLE II: The training cost and accuracy comparison between different approaches, using ResNet as backbone
     Approaches      Relative L2 error      Training cost
     Real      Imag.      (hours)
     Conventional in full out      0.7411      0.7391      11.956
     Conventional in residual out      0.9479      0.9517      11.912
     Our method with full out      0.5730      0.5671      11.968
     Our method with residual out      0.5748      0.5653      11.940
TABLE III: The training cost and accuracy comparison between different approaches, using Unet as backbone
     Approaches      Relative L2 error      Training cost
     Real      Imag.      (hours)
     Conventional in full out      0.3050      0.3074      2.394
     Conventional in residual out      0.4120      0.4160      2.417
     Our method with full out      0.1963      0.1957      2.363
     Our method with residual out      0.1905      0.1892      2.375